OpenStudy (anonymous):

Can someone please tell me if I am correct? The distance between the end of a lamppost's shadow and the top of the lamppost is 39 feet. The angle measure of the shadow is 23°. How long is the shadow of the lamppost, rounded to the nearest tenth of a foot? Type the numeric answer only in the box below. Numerical Answers Expected!

OpenStudy (anonymous):

Is the correct answer 36?

OpenStudy (ivettef365):

cos23 = x/39 therefore 39 cos23

OpenStudy (anonymous):

cos23=x/39 x=39*cos23 x=36

OpenStudy (ivettef365):

Good

OpenStudy (anonymous):

Ok thank you both, wanted to make sure I was doing it correct.

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Can someone look at his one? I had 36 but it wasn't right? Please

OpenStudy (ivettef365):

That would be cos23 = x/39

OpenStudy (ivettef365):

It says to the nearest tenth therefore you need to write 35.9

OpenStudy (anonymous):

Oh ok , I mis read it. Thanks